OpenStudy (aaronandyson):

If 2sin(theta) = 2 - cos theta ,find sin theta

OpenStudy (agent0smith):

You could square both sides. First get the cosine alone, that way you can use the identity. \[\large 2\sin \theta -2 = - \cos \theta \]then square both sides \[\large (2\sin \theta -2)^2 = \cos^2 \theta \] then you can replace cos^2 theta. \[\large (2\sin \theta -2)^2 = 1- \sin^2 \theta\]Now expand the left, and get it equal to zero, and it should be possible to factor like a quadratic.

OpenStudy (aaronandyson):

2sin^2 theta + 4 - 2(2sin(theta)(2)

OpenStudy (agent0smith):

Check your math.

OpenStudy (aaronandyson):

4sin^2 theta + 4 - 2(2sin(theta)(2)

OpenStudy (agent0smith):

idk what 2(2sin(theta)(2) is meant to be but clean it up.

OpenStudy (aaronandyson):

8 sin theta?

OpenStudy (agent0smith):

Then write 8 sin theta...

OpenStudy (agent0smith):

And you're missing an equals sign. \[\large 4\sin^2 \theta -8 \sin \theta +4 = 1- \sin^2 \theta\]

OpenStudy (aaronandyson):

Do I have to move the - sin^2 theta from the RHS to the LHS

OpenStudy (agent0smith):

Yes, get it equal to zero.

OpenStudy (aaronandyson):

\[5 \sin^2 \theta - 8 \sin \theta + 3 = 0\]

OpenStudy (agent0smith):

Now factor as if it was 5x^2 - 8x +3 = 0

OpenStudy (aaronandyson):

-5 and 3 ?

OpenStudy (aaronandyson):

5x^2 - 5x - 3x +3 = 0

OpenStudy (aaronandyson):

5x(x-1)-3(x-1) (5x-3)(x-1)

OpenStudy (aaronandyson):

x = 3/5 ot x = 1

OpenStudy (aaronandyson):

or*

OpenStudy (agent0smith):

Don't forget it's not actually 5x^2 - 8x +3 = 0...

OpenStudy (aaronandyson):

I do know that. So now theta = 3/5or 1 ,right?

OpenStudy (aaronandyson):

or sin theta = 1 theta = 90

OpenStudy (agent0smith):

Look at the question again, you aren't asked for theta, you're asked for sine theta.

OpenStudy (aaronandyson):

3/5 and 1 can be the values of sin theta

OpenStudy (agent0smith):

Remember you replaced x with sin theta, so de-replace and you have sin theta = 3/5 and sin theta = 1